# TestART: Improving LLM-based Unit Testing via Co-evolution of Automated Generation and Repair Iteration

### v5

@@
Abstract
Unit testing is crucial for detecting bugs in individual program units but consumes time and effort. Recently, large language models (LLMs) have demonstrated remarkable capabilities in generating unit test cases. However, several problems limit their ability to generate high-quality unit test cases: (1) compilation and runtime errors caused by the hallucination of LLMs; (2) lack of testing and coverage feedback information restricting the increase of code coverage;(3) the repetitive suppression problem causing invalid LLM-based repair and generation attempts. To address these limitations, we propose TestART, a novel unit test generation method. TestART improves LLM-based unit testing via co-evolution of automated generation and repair iteration, representing a significant advancement in automated unit test generation. TestART leverages the template-based repair strategy to effectively fix bugs in LLM-generated test cases for the first time. Meanwhile, TestART extracts coverage information from successful test cases and uses it as coverage-guided testing feedback. It also incorporates positive prompt injection to prevent repetition suppression, thereby enhancing the sufficiency of the final test case. This synergy between generation and repair elevates the correctness and sufficiency of the produced test cases significantly beyond previous methods. In comparative experiments, TestART demonstrates an 18% improvement in pass rate and a 20% enhancement in coverage across three types of datasets compared to baseline models. Additionally, it achieves better coverage rates than EvoSuite with only half the number of test cases. These results demonstrate TestART’s superior ability to produce high-quality unit test cases by harnessing the power of LLMs while overcoming their inherent flaws.
Index Terms:
@@ Abstract
1 Introduction
Unit testing is crucial in software development and is the basic procedure of the software testing process. The primary goal of unit testing is to generate a suite of unit test cases that can improve the coverage and detect errors of the targeted software early in the development process. In unit testing, the focal method, which generally refers to the specific method or function being tested, is the primary target of most unit tests. However, manually creating and maintaining unit test cases can be laborious and time-consuming [1]. To mitigate these difficulties, researchers propose various methodologies and tools to automate the unit test generation process.
@@ 1 Introduction
Based on different technological foundations and application scenarios, the automated unit test generation tools can be categorized into two main types [2]: traditional program analysis-based [3, 4] and language-model-based [5, 1, 6, 7, 8]. Traditional program analysis-based methods include search-based [9, 10, 11, 12], constraint-based [13, 14], and random-based [15, 16] techniques. However, the most widely used search-based software testing (SBST) tools (e.g., EvoSuite [3] and Pynguin [4]) generate test cases that differ significantly from human-written ones. This makes it challenging to read, understand, or modify to reuse the generated unit test cases [17]. Language model-based approaches based on transformer architecture [18] like A3Test [5] and AthenaTest [1] can learn from real-world focal methods and generate developer-written test cases. However, they cannot even correct a syntax error by engaging with the model, resulting in most generated tests being incorrect and low passing rates. Recently, large language models (LLMs, e.g. ChatGPT [19]) demonstrate improved performance over previous language-model-based methods on unit testing, represented by several works [6, 7, 8, 20]. These methods can generate test cases that approximate human-level comprehension and deliver relatively high coverage, making them a central focus in this field.
@@
However, these methods remain constrained by the instability and hallucination issues inherent in LLMs. While fine-tuning techniques have been proposed to mitigate hallucination problems, they tend to increase both time and financial costs. Additionally, ineffective prompt engineering, along with an insufficient understanding of environmental dependencies, can lead to compilation errors due to invalid context information. Furthermore, LLMs can barely run and analyze the generated test cases [6], so they are hard to obtain the testing feedback, including error report or coverage feedback [2], which may cause runtime errors and low coverage rates. Although developers can re-enter the feedback information from compilers to LLMs and request them to repair bugs, the inevitable faithfulness hallucination [21] problem could make this self-repair interaction into potentially endless iterations. Even with continuous attempts, the test cases remain difficult to optimize due to the impact of the repetitive suppression problem [22, 23].
@@
To address the aforementioned issues and mitigate the impact of hallucinations, in this paper, we propose TestART, which adequately addresses the above limitation. TestART is a novel unit test generation method enhanced by the co-evolution of automated generation and repair iteration based on the LLMs (e.g., ChatGPT-3.5). We aim to utilize the generation capability of LLMs and the synergy of the generation-repair mechanism to create high-quality unit test cases. TestART starts by pre-processing the source code and then utilizes the LLMs to generate an initial set of test cases. These test cases undergo a loop involving compilation, execution and repair to identify, capture and fix errors based on TestART’s repair strategy. Then, TestART evaluates the coverage rates of executed test cases. If the coverage meets a high standard, the test cases are output as the final result. If not, we use the positive prompt injection and the coverage-guided testing feedback to alleviate the hallucination and incrementally generate test cases for the next iteration. As a result, TestART can automatically generate test cases with high passing execution and coverage rates.
@@
To verify the effectiveness of TestART, we conduct extensive experiments to compare TestART with state-of-the-art automated unit test generation approaches on open-source and industrial benchmarks. The experimental results show that the pass rate of test cases generated by TestART can far exceed other methods, reaching 78.55% on Defects4J, with an improvement of 18% compared to both ChatGPT-4.0 model and ChatUniTest. In addition, TestART achieves an average line coverage rate of 90.96% on passed focal methods, exceeding the SBST tool EvoSuite by 3.4% with only half the number of test cases. We also complete ablation experiments to verify the effectiveness of modularization. The experimental results show that TestART effectively utilizes LLMs’ generation capability and generates high-quality unit test cases. To sum up, the main contributions of this paper are as follows:
@@
Figure 1: The common workflow of LLM-based unit testing generation method
Figure 2: The motivation of TestART
2 Background and Motivation
@@
2.1 LLM-based Unit Test Generation
Nowadays, with the explosion of LLMs represented by ChatGPT, the LLM-based unit test generation methods attract more attention. To reduce time and financial costs, mainstream researchers reduced their reliance on fine-tuning strategies and instead shifted towards exploring viable prompting mechanisms for generating test cases [24, 6, 8, 25]. Based on previous works, Fig. 1 illustrates the commonly used workflow for constructing prompts and performing testing validation to generate test cases using LLMs. Most of the approaches consist of three main modules: prompt constructor, test validator and processing engine. Firstly, the prompt constructor initializes system prompts, defining the role of the LLMs as a “testing expert”. Then, the source code is taken as part of different user prompts, instructing LLMs to generate the test cases. Secondly, after getting the generated test cases, the test validator invokes the compiler and controller to perform dynamic analysis on the test cases. If the test cases do not compile or execute successfully, they will be subject to the processing engine along with the error message. Thirdly, researchers use various engines, including static analysis, error message parser, or LLM-repair, to further process the results and update the user prompt to improve the LLM-generated test cases. Ultimately, after several iterations, these methods yield test cases of comparable quality.
2.2 Motivation
@@
The main insight from this example is that although LLM-generated test cases exhibit high logic quality, they are frequently compromised by simple errors, adversely affecting the pass and coverage rate. In fact, the pass rate of test cases directly generated by ChatGPT does not exceed 50% [8]. Meanwhile, due to the issue of repetitive suppression in LLMs [23], the output from LLMs often includes repetitive or highly similar results, containing similar errors. Therefore, we propose applying the automated program repair (APR) technique that utilizes predefined templates to precisely and reliably correct these errors rather than only relying on LLMs to repair bugs.
Figure 3: The Overview of TestART
3 Approach
In this section, we first present the overview of the proposed TestART, followed by a detailed examination of its key modules. The workflow of TestART is illustrated in Fig. 3. Given a source code, TestART initially performs pre-processing (as described in Sec. 3.1) to alleviate the hallucination problem. Afterward, TestART leverages the ChatGPT-3.5 model to generate the initial set of test cases $T_{candidate}$ (as explained in Sec. 3.2). Once generated, $T_{candidate}$ enters the co-evolution loop (depicted by the yellow arrows in the figure), which aims to produce high-quality final test cases. Specifically, TestART submits $T_{candidate}$ for compilation and execution to capture error messages. Based on the captured errors (compilation errors, runtime errors, or other detected bugs), TestART applies the repair strategy to address these errors and recompiles and runs until it passes without bugs or reaches the maximum number of iterations. If the test cases still unsuccess, $T_{candidate}$ is discarded. Otherwise, $T_{candidate}$ is promoted to $T_{success}$. Next, TestART employs JUnit and OpenClover to evaluate the coverage of $T_{success}$ over the source code, transforming the uncovered areas into test feedback. If $T_{success}$ meets the coverage standard, it will be output as the final result ($T_{final}$). Otherwise, $T_{success}$ and the coverage-guidance feedback are provided as positive prompt injections to the ChatGPT-3.5 model, continuing to the next co-evolution loop and revert to $T_{candidate}$. A detailed description of the co-evolution of automated generation and repair is introduced in Sec. 3.3. Benefiting from this synergy, TestART can iterate test cases incrementally and ensure that each round of test cases passes successfully, which continuously improves coverage.
3.1 Pre-processing
@@ 3.3 Synergy
3.3.1 Repair Strategy
By analyzing the errors in the initial $T_{candidate}$, we find that the LLM-generated test cases exhibit systematic and distinct errors. Moreover, due to the problem of repetitive suppression, these errors tend to recur. Because these errors are inherently produced by the LLMs, correcting them using the models alone proves challenging. Consequently, we developed repair templates based on expert knowledge, drawing upon the most advanced traditional repair techniques. Specifically, TestART first identifies the location of the erroneous code and extracts the error messages. We separate the compilation and execution of the test code into two distinct steps to more accurately identify the source of any errors and apply the corresponding repair templates. For compilation errors, TestART identifies issues by analyzing the logs generated during compilation. These logs contain detailed error messages, the location in the code where the error occurred, and relevant warning information. For runtime errors, TestART locates errors by examining the stack trace information. Then TestART utilizes corresponding templates to repair the test code. To address spelling and range errors that template-based repair cannot cover, TestART attempts a one-time LLM-based repair after multiple failed template repair attempts. Unlike previous repair strategies applied to semantic buggy code [26, 27, 28], TestART focuses on repairing compilation errors (including syntax errors, import errors, and scope errors) and runtime errors that appear in the generated test cases. Our repair strategy focuses more on the internal logic and expected behavior of the test cases, aiming at increasing the correctness. TestART can repair the most common errors in LLM-generated test cases by designing five experiential repair templates, which are simple, stable and effective (reaching 50% and 75% of compile and runtime error repair rates, respectively). In the following, we demonstrate the five templates and how the buggy test code can be repaired using the designed templates.
@@ 3.3.1 Repair Strategy
3.3.2 Coverage-guided Testing Feedback
Figure 4: The prompt of LLM-based repair
@@ 3.3.2 Coverage-guided Testing Feedback
3.3.3 Positive Prompt Injection
Figure 5: The process of positive prompt injection and coverage-guided testing feedback
@@ 3.3.3 Positive Prompt Injection
Prompt injection technique is used for attacking and defending LLMs [29]. It involves inserting unexpected or malicious content into the prompt to change the system’s behavior, extract unauthorized data, or trigger unintended actions. The core is that altering the output of the LLMs induces LLMs to perceive the modified content as its original generation. However, unlike injecting attacks, which are considered negative, TestART treats the repaired test cases $T_{success}$ as positive injections to mitigate hallucination.
@@ 4 Experiment Design
RQ4: What is the performance of TestART when applied to unlearned datasets, and how well does it generalize to new data?
4.1 Datasets
TABLE I: The description of datasets used in the experiment
@@ 4.1 Datasets
Defects4J [30] is a collection of reproducible software bugs and supporting infrastructure to advance research in software engineering. We follow previous work [6] and select five Java projects from Defects4J to evaluate TestART, shown in Table I. We extracted public and non-abstract classes from five projects as focal classes and extracted their public methods as focal methods, a total of 8192 focal methods extracted.
HITS dataset [25] comprises internet-sourced projects developed after the knowledge cutoff date for ChatGPT-3.5, indicating that the LLMs have no training data from this dataset. The HITS dataset contains 885 focal methods from 3 projects. We retain the part of the dataset in the original paper that is not duplicated in Defects4J as this dataset.
Internal dataset is an internal Java unit test dataset provided by Huawei. This dataset contains a total of 668 completely internal Java focal methods. This means that the open-source model cannot learn the data features of this part. The specific configuration of the dataset is shown in the Table I.
4.2 Baselines
To evaluate the effectiveness of our proposal, we compare TestART with four baselines. We choose the state-of-the-art methods of four kinds of solutions as baselines including the SBST tool (EvoSuite), the language-model-based method (A3Test), the Large language model (two ChatGPT models) and the LLM-based test case generation approach (ChatUniTest).
EvoSuite [3] is a traditional unit test generation tool that uses evolutionary algorithms to create new test cases and a fitness function to direct the search process. This approach helps achieve high code coverage criteria, such as branch and line coverage.
A3Test [5] is a test case generation approach based on deep learning that is enhanced with assertion knowledge and includes a mechanism to verify naming consistency and test signatures. A3Test applies domain adaptation principles to adapt existing knowledge from an assertion generation task to the test case generation task.
ChatGPT [19] is a highly advanced technology that can replicate human speech and reasoning by learning from a vast library of human communication. It achieves performance levels comparable to humans in professional and academic settings. ChatGPT-3.5 and ChatGPT-4.0 are set as two different baselines.
ChatUniTest [6] is a ChatGPT-based automated unit test generation tool developed by the Generation-Validation-Repair framework. ChatUniTest generates tests by analyzing the project, extracting critical information, and creating an adaptive context that includes the focal method and its dependencies within a set token limit.
4.3 Evaluation Metrics

### v6

@@
Abstract.
Unit testing is crucial for detecting bugs in individual program units but consumes time and effort. Recently, large language models (LLMs) have demonstrated remarkable capabilities in generating unit test cases. However, several problems limit their ability to generate high-quality unit test cases: (1) compilation and runtime errors caused by the hallucination of LLMs; (2) lack of testing and coverage feedback information restricting the increase of code coverage; and (3) the repetitive suppression problem causing invalid LLM-based repair and generation attempts. To address these limitations, we propose TestART, a novel unit test generation method. TestART improves LLM-based unit testing via co-evolution of automated generation and repair iteration, representing a significant advancement in automated unit test generation. TestART leverages the template-based repair strategy to effectively fix bugs in LLM-generated test cases for the first time. Meanwhile, TestART extracts coverage information from successful test cases and uses it as coverage-guided testing feedback. It also incorporates positive prompt injection to prevent repetition suppression, thereby enhancing the sufficiency of the final test case. This synergy between generation and repair elevates the correctness and sufficiency of the produced test cases significantly beyond previous methods. Through comparative experiments, TestART demonstrates an 18% improvement in pass rate and a 20% enhancement in coverage across three types of datasets compared to baselines. Additionally, it achieves better coverage rates than EvoSuite with only half the number of test cases. These results demonstrate TestART’s superior ability to produce high-quality unit test cases by harnessing the power of LLMs while overcoming their inherent flaws.
@@ Abstract.
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: TOSEMjournalvolume: 0journalnumber: 0article: 1publicationmonth: 0ccs: Software and its engineering Software testing and debuggingccs: Computing methodologies Natural language processing
1. Introduction
@@
Unit testing is crucial in software development and is the basic procedure of the software testing process. The primary goal of unit testing is to generate a suite of unit test cases that can improve code coverage and detect errors in the targeted software early during the development process. In unit testing, the focal method, which generally refers to the specific method or function being tested, is the primary target of most unit tests. However, manually creating and maintaining unit test cases can be laborious and time-consuming (Tufano et al., 2020a). To mitigate these difficulties, researchers propose various methodologies and tools to automate the unit test generation process. Based on different technological foundations and application scenarios, the automated unit test generation tools can be categorized into two main types (Tang et al., 2024): traditional program analysis-based (Fraser and Arcuri, 2011; Lukasczyk and Fraser, 2022) and language-model-based (Alagarsamy et al., 2024a; Tufano et al., 2020a; Chen et al., 2024; Bhatia et al., 2024; Yuan et al., 2024). Traditional program analysis-based methods include search-based (Baresi and Miraz, 2010; Derakhshanfar et al., 2022; Fraser and Zeller, 2010; Harman and Jones, 2001), constraint-based (Ma et al., 2015; Sakti et al., 2014), and random-based (Andrews et al., 2011; Pacheco et al., 2007) techniques. However, the most widely used search-based software testing (SBST) tools (e.g., EvoSuite (Fraser and Arcuri, 2011) and Pynguin (Lukasczyk and Fraser, 2022)) generate test cases that differ significantly from human-written ones. This makes it challenging to read, understand, or modify to reuse the generated unit test cases (Gargari and Keyvanpour, 2021). Language model-based approaches based on transformer architecture (Vaswani et al., 2017) like A3Test (Alagarsamy et al., 2024a) and AthenaTest (Tufano et al., 2020a) can learn from real-world focal methods and generate developer-written test cases. However, they cannot even correct a syntax error by engaging with the model, resulting in most generated tests being incorrect and low passing rates. Recently, large language models (LLMs, e.g., ChatGPT (OpenAI, 2023)) demonstrate improved performance over previous language-model-based methods on unit testing, represented by several works (Chen et al., 2024; Bhatia et al., 2024; Yuan et al., 2024; Alshahwan et al., 2024). These methods can generate test cases that approximate human-level comprehension and deliver relatively high coverage, making them a central focus in this field.
@@
However, these methods remain constrained by the instability and hallucination issues inherent in LLMs. While fine-tuning techniques have been proposed to mitigate hallucination problems, they tend to increase both time and financial costs. Additionally, ineffective prompt engineering, along with an insufficient understanding of environmental dependencies, can lead to compilation errors due to invalid context information. Furthermore, LLMs can barely run and analyze the generated test cases (Chen et al., 2024), so they are hard to obtain the testing feedback, including error report or coverage feedback (Tang et al., 2024), which may cause runtime errors and low coverage rates. Although developers can re-enter the feedback information from compilers to LLMs and request them to repair bugs, the inevitable faithfulness hallucination (Huang et al., 2023) problem could make this self-repair interaction into potentially endless iterations. Even with continuous attempts, the test cases remain difficult to optimize due to the impact of the repetitive suppression problem (Xu et al., 2022; Zhang et al., 2023c).
@@
To address the aforementioned issues and mitigate the impact of hallucinations, in this paper, we propose TestART. TestART is a novel unit test generation method enhanced by the co-evolution of automated generation and repair iteration based on the LLMs (e.g., ChatGPT-3.5). We aim to utilize the generation capability of LLMs and the synergy of the generation-repair mechanism to create high-quality unit test cases. TestART starts by pre-processing the source code and then utilizes the LLMs to generate an initial set of test cases. These test cases undergo a loop involving compilation, execution and repair to identify, capture and fix errors based on TestART’s repair strategy. TestART uses a template-driven repair strategy tailored to the outputs of LLMs, prioritizing template-based repairs and resorting to LLM-based repair only for unresolved cases. After that, TestART evaluates the coverage rates of executed test cases. If the coverage meets a high standard, the test cases are output as the final result. If not, we use the positive prompt injection and the coverage-guided testing feedback to alleviate the hallucination and incrementally generate test cases for the next iteration. As a result, TestART can automatically generate test cases with high passing execution and coverage rates under the co-evolution between the generation and repair processes, reducing manual intervention significantly.
@@
To verify the effectiveness of TestART, we conduct extensive experiments to compare TestART with state-of-the-art automated unit test generation approaches on open-source and industrial benchmarks. The experimental results show that the pass rate of test cases generated by TestART can far exceed other methods, reaching 78.55% on Defects4J, with an improvement of 18% compared to both ChatGPT-4.0 model and ChatUniTest. In addition, TestART achieves an average line coverage rate of 90.96% on passed focal methods, exceeding the SBST tool EvoSuite by 3.4% with only half the number of test cases. We also complete ablation experiments to verify the effectiveness of modularization. The experimental results show that TestART effectively utilizes LLMs’ generation capability and generates high-quality unit test cases.
To sum up, the main contributions of this paper are as follows:
@@
2. Background and Motivation
Figure 1. The common workflow of LLM-based unit testing generation method
Figure 2. The motivation of TestART
@@ 2. Background and Motivation
2.1. LLM-based Unit Test Generation
Nowadays, with the explosion of LLMs represented by ChatGPT, the LLM-based unit test generation methods attract more attention. To reduce time and financial costs, mainstream researchers reduced their reliance on fine-tuning strategies and instead shifted towards exploring viable prompting mechanisms for generating test cases (Schäfer et al., 2023; Chen et al., 2024; Yuan et al., 2024; Wang et al., 2024b). Based on previous works, Fig. 1 illustrates the commonly used workflow for constructing prompts and performing testing validation to generate test cases using LLMs. Most of the approaches consist of three main modules: prompt constructor, test validator and processing engine. Firstly, the prompt constructor initializes system prompts, defining the role of the LLMs as a “testing expert”. Then, the source code is taken as part of different user prompts, instructing LLMs to generate the test cases. Secondly, after getting the generated test cases, the test validator invokes the compiler and controller to perform dynamic analysis on the test cases. If the test cases do not compile or execute successfully, they will be subject to the processing engine along with the error message. Thirdly, researchers use various engines, including static analysis, error message parser, or LLM-repair, to further process the results and update the user prompt to improve the LLM-generated test cases. Ultimately, after several iterations, these methods yield test cases of comparable quality.
2.2. Motivation
@@
The main insight from this example is that although LLM-generated test cases exhibit high logic quality, they are frequently compromised by simple errors, adversely affecting the pass and coverage rate. In fact, the pass rate of test cases directly generated by ChatGPT does not exceed 50% (Yuan et al., 2024). Meanwhile, due to the issue of repetitive suppression in LLMs (Zhang et al., 2023c), the output from LLMs often includes repetitive or highly similar results, containing similar errors. Therefore, we propose applying the automated program repair (APR) technique that utilizes predefined templates to precisely and reliably correct these errors rather than only relying on LLMs to repair bugs.
3. Approach
Figure 3. The Overview of TestART
In this section, we first present the overview of the proposed TestART, followed by a detailed examination of its key modules. The workflow of TestART is illustrated in Fig. 3. Given a source code, TestART initially performs pre-processing (as described in Sec. 3.1) to alleviate the hallucination problem. Afterward, TestART leverages the ChatGPT-3.5 model to generate the initial set of test cases $T_{candidate}$ (as explained in Sec. 3.2). Once generated, $T_{candidate}$ enters the co-evolution loop (depicted by the yellow arrows in Fig. 3), which aims to produce high-quality final test cases. Specifically, TestART submits $T_{candidate}$ for compilation and execution to capture error messages. Based on the captured errors (compilation errors, runtime errors, or other detected bugs), TestART applies the repair strategy to address these errors and recompiles and runs until it passes without bugs or reaches the maximum number of iterations. If the test cases still unsuccess, $T_{candidate}$ is discarded. Otherwise, $T_{candidate}$ is promoted to $T_{success}$. Next, TestART employs JUnit and OpenClover to evaluate the coverage of $T_{success}$ over the source code, transforming the uncovered areas into test feedback. If $T_{success}$ meets the coverage standard, it will be output as the final result ($T_{final}$). Otherwise, $T_{success}$ and the coverage-guidance feedback are provided as positive prompt injections to the ChatGPT-3.5 model, continuing to the next co-evolution loop and revert to $T_{candidate}$. A detailed description of the co-evolution of automated generation and repair is introduced in Sec. 3.3. Benefiting from this synergy, TestART can iterate test cases incrementally and ensure that each round of test cases passes successfully, which continuously improves coverage.
3.1. Pre-processing
@@ 3.3. Synergy
3.3.1. Repair Strategy
By analyzing the errors in the initial $T_{candidate}$, we find that the LLM-generated test cases exhibit systematic and distinct errors. Moreover, due to the problem of repetitive suppression, these errors tend to recur. Because these errors are inherently produced by the LLMs, correcting them using the models alone proves challenging. Consequently, we developed repair templates based on expert knowledge, drawing upon the most advanced traditional repair techniques. Specifically, TestART first identifies the location of the erroneous code and extracts the error messages. We separate the compilation and execution of the test code into two distinct steps to more accurately identify the source of any errors and apply the corresponding repair templates. For compilation errors, TestART identifies issues by analyzing the logs generated during compilation. These logs contain detailed error messages, the location in the code where the error occurred, and relevant warning information. For runtime errors, TestART locates errors by examining the stack trace information. Then TestART utilizes corresponding templates to repair the test code. To address spelling and range errors that template-based repair cannot cover, TestART attempts a one-time LLM-based repair after multiple failed template repair attempts. Unlike previous repair strategies applied to semantic buggy code (Liu et al., 2019b; Zhang et al., 2023b; Liu et al., 2019a), TestART focuses on repairing compilation errors (including syntax errors, import errors, and scope errors) and runtime errors that appear in the generated test cases. Our repair strategy focuses more on the internal logic and expected behavior of the test cases, aiming at increasing the correctness. TestART can repair the most common errors in LLM-generated test cases by designing five experiential repair templates, which are simple, stable and effective (reaching 50% and 75% of compile and runtime error repair rates, respectively). In the following, we demonstrate the five templates and how the buggy test code can be repaired using the designed templates.
@@ 3.3.1. Repair Strategy
3.3.2. Coverage-guided Testing Feedback
Figure 4. The prompt of LLM-based repair
@@ 3.3.2. Coverage-guided Testing Feedback
3.3.3. Positive Prompt Injection
Figure 5. The process of positive prompt injection and coverage-guided testing feedback
@@ 3.3.3. Positive Prompt Injection
Prompt injection technique is used for attacking and defending LLMs (Choi et al., 2022). It involves inserting unexpected or malicious content into the prompt to change the system’s behavior, extract unauthorized data, or trigger unintended actions. The core is that altering the output of the LLMs induces LLMs to perceive the modified content as its original generation. However, unlike injecting attacks, which are considered negative, TestART treats the repaired test cases $T_{success}$ as positive injections to mitigate hallucination.
@@ 4. Experiment Design
RQ4: What is the performance of TestART when applied to unlearned datasets? and how well does it generalize to new data?
RQ5: What is the time and money cost of TestART to generate test cases?
4.1. Datasets
Table 1. The description of datasets used in the experiment
@@ 4.1. Datasets
Defects4J (Just et al., 2014) is a collection of reproducible software bugs and supporting infrastructure to advance research in software engineering. We follow previous work (Chen et al., 2024) and select five Java projects from Defects4J to evaluate TestART, shown in Table 1. We extracted public and non-abstract classes from five projects as focal classes and extracted their public methods as focal methods, a total of 8192 focal methods extracted.
HITS dataset (Wang et al., 2024b) comprises internet-sourced projects developed after the knowledge cutoff date for ChatGPT-3.5, indicating that the LLMs have no training data from this dataset. The HITS dataset contains 885 focal methods from 3 projects. We retain the part of the dataset in the original paper that is not duplicated in Defects4J as this dataset.
Internal dataset is an internal Java unit test dataset provided by Huawei. This dataset contains a total of 668 completely internal Java focal methods. This means that the open-source model cannot learn the data features of this part. The specific configuration of the dataset is shown in the Table 1.
4.2. Baselines
To evaluate the effectiveness of our proposal, we compare TestART with four baselines. We choose the state-of-the-art methods of four kinds of solutions as baselines including the SBST tool (EvoSuite), the language-model-based method (A3Test), the Large language model (two ChatGPT models) and the LLM-based test case generation approach (ChatUniTest). Note that CHATTESTER is excluded from baseline comparisons due to 1) the similar LLM-based repair methodology with ChatUnitTest and 2) ChatUnitTest shows better performance than CHATTESTER in the experiments according to HITS(Wang et al., 2024b).
EvoSuite (Fraser and Arcuri, 2011) is a traditional unit test generation tool that uses evolutionary algorithms to create new test cases and a fitness function to direct the search process. This approach helps achieve high code coverage criteria, such as branch and line coverage.
A3Test (Alagarsamy et al., 2024a) is a test case generation approach based on deep learning that is enhanced with assertion knowledge and includes a mechanism to verify naming consistency and test signatures. A3Test applies domain adaptation principles to adapt existing knowledge from an assertion generation task to the test case generation task.
ChatGPT (OpenAI, 2023) is a highly advanced technology that can replicate human speech and reasoning by learning from a vast library of human communication. It achieves performance levels comparable to humans in professional and academic settings. ChatGPT-3.5 and ChatGPT-4.0 are set as two different baselines.
ChatUniTest (Chen et al., 2024) is a ChatGPT-based automated unit test generation tool developed by the Generation-Validation-Repair framework. ChatUniTest generates tests by analyzing the project, extracting critical information, and creating an adaptive context that includes the focal method and its dependencies within a set token limit.
4.3. Evaluation Metrics